# Diversity-Enhanced Reasoning for Subjective Questions

### v1

@@ E.1 Formulation of the Diversity Metric
Ours 73.77 92.22 70.51 57.20 95.56 42.94 72.69 81.99 68.59 555.0 74.87
Ours(+SFT+GRPO) 80.98 89.36 60.29 63.31 96.36 16.85 65.16 93.74 73.35 620.0 82.16
@@
Table 9: Detailed composition of the diversity scores based on the output of R1-Distilled-Qwen-7B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
@@
Ours(+SFT) 85.48 90.17 61.54 59.99 97.89 40.09 69.97 94.78 72.11 1364 81.36
Ours(+SFT+GRPO) 94.56 88.19 61.25 59.31 97.83 48.44 63.77 95.48 75.80 1551.4 87.89
@@
Table 10: Detailed composition of the diversity scores based on the output of R1-Distilled-Llama-8B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
@@
Table 11: Detailed composition of the diversity scores based on the output of R1-Distilled-Qwen-14B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and normalized diversity score.
@@
Table 12: Detailed composition of the diversity scores based on the output of Qwen3-8B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
E.2 Embedding Similarity as the Diversity Metric
$RODS=\frac{2}{n(n-1)}\sum_{i<j}d_{ij}$

### v3

@@ F.2 Embedding Similarity as the Diversity Metric
Ours(+SFT) 73.77 92.22 70.51 57.20 95.56 42.94 72.69 81.99 68.59 620.0 74.87
Ours(+SFT+GRPO) 80.98 89.36 60.29 63.31 96.36 16.85 65.16 93.74 73.35 555.0 82.16
@@
Table 10: Detailed composition of the diversity scores based on the output of R1-Distilled-Llama-8B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. In addition, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
@@
Ours(+SFT) 85.48 90.17 61.54 59.99 97.89 40.09 69.97 94.78 72.11 1551.4 81.36
Ours(+SFT+GRPO) 94.56 88.19 61.25 59.31 97.83 48.44 63.77 95.48 75.80 1364 87.89
@@
Table 11: Detailed composition of the diversity scores based on the output of R1-Distilled-Qwen-14B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. In addition, we also provide the combined diversity score, average reasoning length and normalized diversity score.
@@
Table 12: Detailed composition of the diversity scores based on the output of Qwen3-8B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. In addition, we also provide the combined diversity score, average reasoning length and length normalized diversity score.